OpenStudy (anonymous):

True or False: When dividing monomials, the monomial with the smaller exponent is cancelled completely.

OpenStudy (anonymous):

True. For example x^5 / x^2 = x^3. The bottom one is canceled completely. Remember, monomials are ones with nothing but a single power of the variable (or a constant).

OpenStudy (anonymous):

This is the case. You need to take it in a general case. I think you are referencing polynomials. Monomial would be x, whereas a polynomial is x^2, x^3,...etc If the numerator is larger, then it will cancel. This follows from laws of logarithms. log(a/b) = a - b So if we have x^4/x^2 then this becomes x^4-2 = x^2 and the bottom goes away. Hope this helps. If it does, please choose Good Answer. Please Like Tutor Sean on facebook. It will help me help others, and I offer free help there with a community of other college tutors. Also check out http://www.tutorsean.net for free videos on math and related subjects.